OpenStudy (camerondoherty):

I'm a bit confused, I have no idea where to start ;-; These are proofs btw. I need to prove if it is true/false \[\csc^2x-1=\cot^2x\] I think I need to start with changing csc to 1/y Then change the 1 to \[\sin^2x + cos^2x\] because it equals one? Some clarification would be nice, thanx c:

OpenStudy (s4sensitiveandshy):

how would you write csc^2x in terms of sin , cos ??

OpenStudy (camerondoherty):

\[\frac{ 1 }{ \sin }^2x\]

OpenStudy (s4sensitiveandshy):

hmm the 2 is flying away :P

OpenStudy (camerondoherty):

Haha, I'll rewrite it \[(\frac{ 1 }{ \sin })^2x\]

OpenStudy (s4sensitiveandshy):

\[\frac{ 1 }{ \sin^2x } -1\] good simplify rewrite it as a single fraction ( find the common denominator )

OpenStudy (camerondoherty):

That's how you rewrite it?

OpenStudy (camerondoherty):

Ooooh, ok

OpenStudy (camerondoherty):

I get it

OpenStudy (camerondoherty):

So \[-\frac{ \sin }{ \sin }\]

OpenStudy (s4sensitiveandshy):

yes because of the fact csc^2x=1/sin^2x i can replace csc^2x with 1/sin^2x

OpenStudy (camerondoherty):

wait \[\frac{ \sin^2x }{ \sin^2x }\]

OpenStudy (s4sensitiveandshy):

hmm no lets say sinx= x \[\frac{ 1 }{ x^2 }-1\] how would you solve this algebra expression ?? :)

jhonyy9 (jhonyy9):

cameron just a moment do you know that cosec x = 1/ sin x yes ?

OpenStudy (camerondoherty):

Yes

jhonyy9 (jhonyy9):

so than cosec^2 x = (1/sin x)^2 yes ?

OpenStudy (camerondoherty):

Yea

jhonyy9 (jhonyy9):

but (1/sin x)^2 = 1/sin^2 x

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (camerondoherty):

Yea, because when you square 1 it stays the same

jhonyy9 (jhonyy9):

exactly

OpenStudy (camerondoherty):

Then you can turn 1 into sin^2x + cos^2x = 1

OpenStudy (camerondoherty):

Cancel out the sines and you have cos

OpenStudy (camerondoherty):

Right?

OpenStudy (s4sensitiveandshy):

let me rewrite it \[\frac{ 1 }{ \sin^2x } -\frac{1}{1}= \cot^2x\] we are working on the left side to find the common denominator multiply top and bottom of the fractions by sin^2x

jhonyy9 (jhonyy9):

sensitive i have one different idea without common denominator

jhonyy9 (jhonyy9):

rewrite the 1 from numerator in the form of sin^2 x +cos^2 x

OpenStudy (s4sensitiveandshy):

\[\frac{ 1 }{ \sin^2x} *\frac{\sin^2x}{\sin^2x} - \frac{1}{1}*\frac{\sin^2x}{\sin^2x}=cot^2x\] \[\frac{ 1 -\sin^2x}{ \sin^2x } =cot^2x\] let me know if you have a question about this ??

OpenStudy (s4sensitiveandshy):

yes thats the easy one btw cameron there are multiple ways to prove trig function :)

OpenStudy (s4sensitiveandshy):

gtg sorry

OpenStudy (camerondoherty):

How did you get the multiplication of sin^2x/sin^2x?

OpenStudy (camerondoherty):

Lol, thanks for helpng though :)

jhonyy9 (jhonyy9):

snsitive the first column of sin^2x/sin^2x not need you write there this is wrong

jhonyy9 (jhonyy9):

so bc. than will get the denominator sin^4 x

jhonyy9 (jhonyy9):

cameron do you see it about what i talk ?

OpenStudy (camerondoherty):

I think I get it now You start with changing csc^2x to 1/sin^2x Then you change the -1 to sin^2x/sin^2x Then You change the cot to cos^2x/sin^2x Since everything is common denominator, you can simplify Right? c:

jhonyy9 (jhonyy9):

no the right part need remain so and from the left part you need getting the right part

OpenStudy (camerondoherty):

Change the 1 to sin^2x + cos^2x?

jhonyy9 (jhonyy9):

first the 1 rewrite in the form of sin^2 /sin^2 x and after this rewrite the 1 from numerator of sin^2 x in this form

jhonyy9 (jhonyy9):

cameron do you understand it now ?

OpenStudy (camerondoherty):

I got it! \[\frac{ 1 }{ \sin^2x }-\frac{ \sin^2x }{ \sin^2x }=\cot^2x\] \[\frac{ 1-\sin^2x }{ \sin^2x }\rightarrow \frac{ \sin^2x+\cos^2x-\sin^2x }{ \sin^2x }\rightarrow \frac{ \cos^2x }{ \sin^2x }=\cot^2x\] Right? c:

jhonyy9 (jhonyy9):

yes this is the right easy way how i thought it hope helped

OpenStudy (camerondoherty):

Thank you so much <3 I didn't understand it, thank you for helping me so c;

jhonyy9 (jhonyy9):

was my pleasure what you dont understand there ? - please ask me now

OpenStudy (camerondoherty):

I get everything c: Thank you for asking

jhonyy9 (jhonyy9):

ok. so good luck bye bye

OpenStudy (s4sensitiveandshy):

i wrote it correctly that's how she got the correct answer:) \[\frac{ 1 }{ \sin^2x } *\frac{\sin^2x}{\sin^2x} - \frac{1}{1} *\frac{\sin^2x}{\sin^2x}\] \[\frac{ 1 }{ \cancel{\sin^2x} } *\frac{\cancel{\sin^2x}}{\sin^2x} - \frac{1}{1} *\frac{\sin^2x}{\sin^2x}\] we will get \[\frac{ 1 }{ \sin^2x } -\frac{\sin^2x}{\sin^2x}\] now you can rewrite as a single fraction\[\frac{ 1-\sin^2x }{ \sin^2x }\] i like how you replaced 1 with sin^2x+cos^2x or you can solve this equation for cos^2x \[\sin^2x+\cos^2x=1 \] \[\cos^2x=1-\sin^2x\] replace the numerator by cos^2x \[\frac{ 1-\sin^2x }{ \sin^2x }=\frac{ cos^2x }{ \sin^2x }\] which is equal to cot^2x